OpenStudy (anonymous):

(2x + y = 6) (-4x + 2y = 4) Solving using elimination?

OpenStudy (anonymous):

2x+y=6 (multiply this equation by 2) -4x+2y=4 4x+2y=12 -4x+2y=4 4y=16 y=4 2x+4=6 2x=2 x=1